# Zero-Shot Scene Change Detection

### v1

@@ 4.2 Experimental Results
Table 1: Experimental results. Results are expressed in per-class IoU and mIoU scores. Despite the absence of a training process, our model outperformed the baseline’s in-domain performance in two out of three subsets.
@@
       Method       Trained Set       Normal         Dust         Dark
@@
Table 2: Experimental results on cross-domain. Results are expressed in the mIoU score across all change classes. We trained the baseline model on each subset and tested it across all subsets. The experimental results show that the baseline model achieves the highest performance when the training set and test set are the same, while performance degrades when the training and test sets differ. In contrast, our method is free from this issue.
Table 1 presents the experimental results with other state-of-the-art, C-3PO [16] from ChangeSim. C-3PO is reproduced under the conditions described in Section 4.1. The baseline model is tested on the same datasets as its training dataset, denoted as in-domain. The table shows that our model achieved superior performance in two out of three subsets: normal and dusty-air, with mIoU of 35.8 and 31.6, respectively. However, our model shows lower performance in low-illumination subset, with mIoU of 25.2.
However, traditional train-based approaches are specialized for the style variation on which they were trained, becoming highly vulnerable when the domains differ. To illustrate this, we conducted additional experiments testing models on different domains from the ones they were trained on. Specifically, we tested the baseline model trained on a particular subset of ChangeSim on the other subsets. The experimental results are shown in Table 2. The baseline model, being specialized for their in-domain data, suffers performance drops when the data changes, indicating a lack of generalization. Specifically, when the baseline model trained on the dusty-air subset is tested on the same dataset, it achieves a performance score of 29.7. However, when the model is trained on the normal or low-illumination datasets and tested on the dusty-air dataset, the performance scores are relatively lower, at 27.2 and 27.1, respectively. This performance drop is also observed in other subsets. This is a major weakness of the baseline in performing the task of change detection, as they are vulnerable to changes in the data. Conversely, our approach, not being tailored to any specific domain, avoids the performance drop typically associated with changes in the dataset In other words, our model can be applied to all subsets without the need for retraining each time the environment changes.
We present qualitative results in Figure 5. These results show that our approach effectively performs change detection without training. More qualitative results and ablation studies will be provided in the supplementary material.
@@
Figure 5: Qualitative Results. Our approach successfully performs change detection across various datasets without training.
5 Conclusion
@@
In this paper, we present a novel approach to zero-shot Scene Change Detection (SCD). Our method performs SCD without training by leveraging a tracking model. To adapt the tracking model for SCD, we introduce two training-free components: the style bridging layer and the adaptive content threshold. Our model demonstrates robustness across various environmental changes, showcasing its versatility. We believe our work offers a fresh perspective on SCD and represents a significant step forward in its practical application.
Limitations. While our model benefits from not requiring a training process, it incurs a relatively higher computational cost during inference compared to existing methods. This increased inference cost arises because our model employs two networks. Future work could focus on enhancing speed by utilizing lighter networks or consolidating certain procedures to improve.
Zero-Shot Scene Change Detection
Supplementary Material
@@ Appendix A Implementation Details
A.1 Details on Mask Generation
This section explains how to create a mask for input into the tracking model. We use Segment Anything Model (SAM) [17] to generate the mask proposal. We executed SAM’s automatic mask generation pipeline with the parameters set as shown in the Table 3.
@@ A.1 Details on Mask Generation
Meanwhile, the masks generated by SAM possess two characteristics that make them unsuitable for our task. First, there are too many small masks. Second, a single pixel can belong to multiple masks. Since the change in scene change detection (SCD) occurs at the object level, the mask size should correspond to the object level and not be too small. Additionally, each pixel in the image can only belong to one object, so it must belong to at most one mask. Therefore, we conducted the following post-processing steps:
Through this process, each pixel belongs to at most one mask, and masks that are too small are naturally removed.
@@
3 Overlay the sorted masks in order. This process ensures that if one pixel belongs
to multiple masks, the largest mask among them is selected.
4 For each mask, if the ratio of the area covered by the overlaid mask is more than 50%,
it is merged with the largest mask among the overlaid masks.
This process removes the masks that occluded most of the area through step 3.
5 For the VL-CMU-CD [29] dataset, remove the masks that do not contain any information.
These are the black areas at the edges of the images.
@@ 4 For each mask, if the ratio of the area covered by the overlaid mask is more than 50%,
A.2 Details on Tracking Model
We use DEVA [13] for our tracking model. We used the DEVA structure with only one modification, incorporating style bridging layers (SBL) within the encoder architecture. The first SBL is positioned immediately after the first convolutional layer, while subsequent SBLs are placed after the addition operation within each residual block [24].
DEVA parameters were set as follows.
@@ A.2 Details on Tracking Model
Appendix B Additional Experiments
We show additional experimental results on the VL-CMU-CD [29] and PCD [2] dataset.
B.1 Dataset and metric
VL-CMU-CD [29] is a dataset that includes information on urban street view changes over a long period, encompassing seasonal variations. Following the baseline approach, we performed predictions using $512\times 512$-sized images. As the change class in this dataset is limited to a binary classification of the ‘missing’ class, we used only the ‘missing’ class for our three types of prediction.
PCD [2] is a dataset consisting of panoramic images and includes two subsets: GSV and TSUNAMI. Following the baselines, we performed predictions on reshaped images of size $256\times 1024$. Each data point is classified into binary change or unchanged categories. We consider the detected new, missing, and replaced predictions into a ‘changed’ class for evaluation.
Evaluation Metrics. Following the previous work, we used the F1 score for evaluation.
B.2 Experimental Results
Table 6 presents the experimental results with other state-of-the-art. The baseline model performs well when the training and test datasets are the same, but its performance drops sharply when the datasets change, indicating its sensitivity to the dataset. Since change detection must operate across various seasons and weather conditions in real-world scenarios, robustness to style changes is essential for effective change detection. This sensitivity implies that the baseline model is overfitted to the limited style changes present in the specific dataset. In contrast, our approach remains robust across all style changes.
VL-CMU-CD & PCD
Test set
Method Trained Set VL-CMU-CD    PCD   Average
C-3PO [16] VL-CMU-CD 79.4 11.6 45.5
PCD 24.3 82.4 53.4
Ours - 51.6 56.5 54.0
Table 6: Experimental results on VL-CMU-CD and PCD. Results are expressed in the F1 score. The baseline model performs best when the training and test are identical. However, its performance greatly declines when these sets differ. Conversely, our method does not exhibit this issue.
Appendix C Ablation Experiments
We conducted ablation experiments to show the effectiveness of our approach. All experiments were conducted on the ChangeSim dataset [6].
@@ Appendix C Ablation Experiments
Addressing Content gap and Style gap. We evaluate the effectiveness of the proposed adaptive content threshold (ACT) and style bridging layer (SBL). As shown in Table 7, experimental results indicate that the combined use of ACT and SBL yields the highest average performance. Additionally, these experiments offer interesting observations: 1. SBL is effective when the style of reference and query image differ (i.e., dusty-air and low-illumination subset), but its effectiveness diminishes in subsets with consistent styles (i.e., normal subset). 2. ACT demonstrates its efficacy particularly when the model’s tracking performance is high (e.g., normal subset). In experiments where tracking performance is poor (e.g., low-illumination subset), the addition of ACT leads to a decline in performance.
@@
      SBL       Normal       Dust       Dark       Average
@@
The number of SBL. We experiment to determine the optimal number of SBL required. We progressively add SBL from the early layer of the encoder. As mentioned in the previous section, the initial Style Bridging Layer (SBL) was placed directly after the first convolutional layer. The following SBLs were inserted after the addition operation within each residual block. As shown in Table 8, experimental results indicate that applying SBL to all blocks of the encoder yields the best performance.
@@
The length of Sequence. We conduct experiments under various $T_{max}$ values. As shown in Table 9, our method shows a significant improvement when extended to video compared to before the extension ($T_{max}=1$). However, it is notable that increasing $T_{max}$ does not consistently lead to improved performance; increasing the video length beyond 60 has little to no impact on performance or may even lead to a decline.
@@
$T_{max}$ Content Threshold Normal Dust Dark Average
@@
Table 10: Ablation study on Adaptive Content Threshold.
The Adaptive Content Threshold. To illustrate the necessity of varying the content threshold according to the sequence length, we conducted experiments across three different sequence lengths. The three sequence lengths were 1, 60, and 30, which are the $T_{max}$ for image-level SCD, our standard $T_{max}$ for video, and an intermediate value, respectively. The fixed threshold values were set to 0.05 and 0.4, approximating the values of ACT when $T_{max}=1$ and $T_{max}=60$, respectively.
As shown in Table 10, when the sequence length is 1, a threshold of 0.05 performs the best, while performance is poor at a threshold of 0.4. Conversely, for sequence lengths of 30 and 60, a threshold of 0.05 results in the lowest performance, while higher thresholds improve performance. Furthermore, the results indicate that the ACT consistently achieves the best performance across all sequence lengths. This shows the validity and effectiveness of ACT.
Appendix D Qualitative Results
@@
Figure 6: Qualitative Results on ChangeSim-Normal.
Figure 7: Qualitative Results on ChangeSim-Dark and ChangeSim-Dust.
Figure 8: Qualitative Results on VL-CMU-CD [29].
Figure 9: Qualitative Results on PCD [2].

### v2

@@ Introduction
Table 1: Experimental results on ChangeSim. The results are expressed in per-class IoU and mIoU scores. Despite the absence of a training process, our model outperformed the baseline’s in-domain performance in two out of three subsets.
@@
 Method Trained Set Normal Dusty-air Low-illum.
@@
Table 2: Experimental results on ChangeSim, cross-domain. The results are expressed in the mIoU score across all change classes. We trained the baseline model on each subset and tested it across all subsets. The experimental results show that the baseline model achieves the highest performance when the training set and test set are the same, while performance degrades when the training and test sets differ. In contrast, our method is free from this issue.
VL-CMU-CD & PCD
Test set
Method Trained Set VL-CMU-CD   PCD  Average
C-3PO VL-CMU-CD 79.4 11.6 45.5
PCD 24.3 82.4 53.4
Ours - 51.6 56.5 54.0
Table 3: Experimental results on VL-CMU-CD and PCD. The results are expressed in the F1 score. The baseline model performs best when the training and test are identical. However, its performance significantly declines when these datasets differ. Conversely, our method is robust to changes in the dataset, maintaining performance without the need for retraining whenever the test environment changes.
Experiments
Experimental Setup
In this section, we briefly introduce the datasets, the relevant settings, and the evaluation metrics.
ChangeSim (Park et al. 2021) is a synthetic dataset with an industrial indoor environment. It includes three subsets with varying environmental conditions: normal, low-illumination, and dusty air. The dataset categorizes changes into four types: new, missing, rotated, and replaced. Despite its variety of environmental variations and change classes, most baseline experiments on this dataset have evaluated only the binary change/unchange classification and have predominantly focused on the normal subset, leaving the dataset’s full potential underexplored. Therefore, we chose the state-of-the-art method, C-3PO (Wang, Gao, and Wang 2023), and reproduced the results under the following conditions: using the original image size ($640\times 480$) to fully utilize the rich information; and including all three subsets. Among the four change classes in this dataset, the rotated class, unlike others, involves slight angular changes of the same object rather than complete appearances or disappearances. We considered this as the object remaining static and integrated this class into static for evaluation.
VL-CMU-CD (Alcantarilla et al. 2018) is a dataset that includes information on urban street view changes over a long period, encompassing seasonal variations. Following the baseline approach, we performed predictions using $512\times 512$-sized images. As the change class in this dataset is limited to a binary classification of the ‘missing’ class, we used only the ‘missing’ class for our three types of prediction.
PCD (Sakurada and Okatani 2015) is a dataset consisting of panoramic images and includes two subsets: GSV and TSUNAMI. Following the baselines, we performed predictions on reshaped images of size $256\times 1024$. Each data point is classified into binary change or unchanged categories. We consider the detected new, missing, and replaced predictions into a ‘changed’ class for evaluation.
Evaluation Metrics
Following the previous work, we employ the mean Intersection over Union (mIoU) metric for ChangeSim and F1 score for VL-CMU-CD and PCD datasets.
@@
Figure 5: Qualitative Results. Our approach successfully performs change detection across various datasets without training. For more qualitative results, see the supplementary material.
Experimental Results
Table 1 presents the experimental results with other state-of-the-art, C-3PO (Wang, Gao, and Wang 2023) from ChangeSim. C-3PO is reproduced under the conditions described in the previous section. The baseline model is tested on the same datasets as its training dataset, denoted as in-domain. The table shows that our model achieved superior performance in two out of three subsets: normal and dusty-air, with mIoU of 35.8 and 31.6, respectively. However, our model shows lower performance in low-illumination subset, with mIoU of 25.2.
However, traditional train-based approaches are specialized for the style variation on which they were trained, becoming highly vulnerable when the domains differ. To illustrate this, we conducted additional experiments testing the baseline model on different domains from the ones it was trained on. Specifically, we tested the baseline model trained on a particular subset of ChangeSim on the other subsets, denoted as cross-domain. The experimental results are shown in Table 2. The baseline model, being specialized for their in-domain data, suffers performance drops when the data changes, indicating a lack of generalization. Specifically, when the baseline model trained on the dusty-air subset is tested on the same subset, it achieves a mIoU of 29.7. However, when the model is trained on the normal or low-illumination subsets and tested on the dusty-air subset, the mIoU scores are relatively lower, at 27.2 and 27.1, respectively. This performance drop is also observed in other cross-domain experiments. Conversely, our approach, which is not tailored to any specific style variation within the training dataset, shows robustness to changing the datasets. In other words, our model can be applied to all subsets without the need for retraining each time the environment changes.
We also conducted experiments on two additional real-world datasets, VL-CMU-CD and PCD. The results are summarized in Table 3. These results reveal a pattern consistent with our previous observations: the baseline model performs well when the training and test datasets are identical, but its performance significantly declines when the datasets differ. In this experiment, the performance drop is particularly severe due to the greater stylistic differences between the datasets. Specifically, the baseline model achieves high in-domain performance with F1 scores of 79.4 and 82.4 on the VL-CMU-CD and PCD datasets, respectively. In contrast, its cross-domain performance drops dramatically to 24.3 and 11.6, respectively. This observation suggests that the baseline model is overfitted to the limited style variations present in the training dataset.
Since the change detection method must function across various seasons and weather conditions in real-world scenarios, robustness to style changes is crucial for reliable performance. The most straightforward way to achieve such robustness against style variations is to employ an approach that does not rely on the training set, as supported by these experimental results.
We present qualitative results in Figure 5. The baseline results correspond to in-domain scenarios, where the training and test datasets are identical. In contrast, our predictions are zero-shot results without any training. The qualitative results show that our approach effectively performs change detection without training. In the supplementary material, we provide more qualitative results and intermediate outcomes of the prediction process.
Ablation experiments
We conducted extensive ablation experiments and more analyses in the supplementary materials to validate the efficacy and robustness of our proposed methods. (1) Performance of zero-shot SCD with and without the proposed adaptive content threshold (ACT) and style bridging layer (SBL). (2) Ablation study based on different SBL settings. (3) Ablation study based on different clip lengths. (4) Ablation study based on different ACT settings.
Conclusion
@@
In this paper, we present a novel approach to zero-shot Scene Change Detection (SCD) applicable to both image and video. Our method performs SCD without training by leveraging a tracking model, which inherently performs change detection between consecutive video frames by recognizing common objects and identifying new or missing objects. To adapt the tracking model specifically for SCD, we propose two training-free components: the style bridging layer and the adaptive content threshold. Through extensive experiments on three SCD datasets, our approach demonstrates its versatility by showing its robustness to various environmental changes. We believe that our work offers a fresh perspective on SCD and represents a significant step forward in its practical application.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-01025, Development of core technology for mobile manipulator for 5G edge-based transportation and manipulation)
Supplementary Material of “Zero-Shot Scene Change Detection”
Technical Appendix
@@ Appendix A Implementation Details
Details on Mask Generation
This section explains how to create a mask for input into the tracking model. We use the Segment Anything Model (SAM) (Kirillov et al. 2023) to generate the mask proposal. We executed SAM’s automatic mask generation pipeline with the parameters set as shown in Table 4.
@@
Meanwhile, the masks generated by SAM exhibit two characteristics that make them unsuitable for our task. First, there are too many small masks. Second, a single pixel can be assigned to multiple masks. These characteristics are problematic because scene change detection (SCD) operates at the object level. Specifically, since the changes occur at the object level, the mask size should correspond to the object level and not be too small. Moreover, each pixel in the image should belong to only one object, meaning it must belong to at most one mask. Therefore, we conducted the post-processing steps outlined in Table 5. This process ensures that each pixel belongs to at most one mask, and small masks are naturally removed.
@@
3 Overlay the sorted masks in order. This process ensures that if one pixel belongs to multiple masks, the largest mask among them is selected.
4 For each mask, if the ratio of the area covered by the overlaid mask is more than 50%, it is merged with the largest mask among the overlaid masks. This process removes the masks that occluded most of the area through step 3.
5 For the VL-CMU-CD (Alcantarilla et al. 2018) dataset, remove the masks that do not contain any information. These are the black areas at the edges of the images.
@@
Details on Tracking Model
We use DEVA (Cheng et al. 2023a) for our tracking model. We used the DEVA structure with only one modification, incorporating style bridging layers (SBL) within the encoder architecture. The first SBL is positioned immediately after the first convolutional layer, while subsequent SBLs are placed after the addition operation within each residual block (He et al. 2016). DEVA parameters were set as shown in Table 6.
Notably, the extension to video with $G$ reduces the need to run the segmentation model $F$ for every frame and reduces total computation cost. This is based on the observation that most tracking models exploit the similarity between adjacent frames in a video to reduce computation: instead of performing computationally intensive operations on all frames, these operations are restricted to keyframes, while intermediate frames are processed through lightweight feature propagation (Zhu et al. 2017; Jain, Wang, and Gonzalez 2019).
In our experiments, $F$ is executed every 5 frames to get the set of object masks (denoted as detection_every in Table 6), while other frames are propagated by $G$ from the previous frame.
@@
Details on Adaptive Content Threshold
The set of missing object masks $M^{t}_{missing}$ defined by equation 7 in the main paper, are the masks that exist in the reference frame but are missing in the query sequence. If an object mask is detected in any frame of the query sequence, it is not classified as missing. Notably, even if an object mask is absent in the majority of frames within the query sequence, the detection of the object mask in just a single frame prevents it from being classified as missing. As a result, as the number of sequences being compared increases, the model becomes more vulnerable to noise, thereby complicating the accurate prediction of missing objects.
Therefore, we hypothesized that the content threshold should adapt to the length of the video sequence. For the reasons mentioned above, a high content threshold risks missing too many true positives in short sequences, while a low content threshold becomes vulnerable to noise in long sequences. To resolve this trade-off, we designed the adaptive content threshold as a parameter explicitly dependent on the video sequence length.
Specifically, we designed the adaptive content threshold to satisfy the following conditions: (1) It should increase as the clip length gets longer. (2) It should have a lower bound to ensure functionality at the image level. (3) It should have an upper bound to prevent excessive growth. Based on these considerations, we derived a simple equation for the adaptive content threshold that depends solely on the sequence length as follows:
$\tau=0.5-{\frac{0.9}{\sqrt{length}+1}}.$ (8)
The effectiveness of this adaptive content threshold is further examined in our ablation studies.
Computing Infrastructure
The experiments were conducted using the Intel Xeon Gold 6426Y CPU and a single NVIDIA RTX A5000 GPU. The software stack employed includes PyTorch 2.1.2 with CUDA version 12.1. The size of memory usage is below 24 GB.
Appendix B Ablation Experiments
We conducted ablation experiments to show the effectiveness of our approach. All experiments were conducted on the ChangeSim dataset (Park et al. 2021). To conserve space in the tables, the names of each subset are abbreviated: dusty-air is abbreviated as ‘Dust’, and low-illumination as ‘Dark’.
Addressing Content gap and Style gap
We evaluate the effectiveness of the proposed adaptive content threshold (ACT) and style bridging layer (SBL). As shown in Table 7, experimental results indicate that the combined use of ACT and SBL yields the highest average performance. Additionally, these experiments offer interesting observations: (1) SBL is effective when the style of reference and query image differ (i.e., dusty-air and low-illumination subset), but its effectiveness diminishes in subsets with consistent styles (i.e., normal subset). (2) ACT demonstrates its efficacy particularly when the model’s tracking performance is high (e.g., normal subset). In experiments where tracking performance is poor (e.g., low-illumination subset), the addition of ACT leads to a decline in performance.
@@ Appendix B Ablation Experiments
The number of SBL
We experiment to determine the optimal number of SBL required. We progressively add SBL from the early layer of the encoder. The first Style Bridging Layer (SBL) was placed directly after the first convolutional layer, and the following SBLs were inserted after the addition operation within each residual block. The experimental results are shown in Table 8. The experimental results indicate that applying SBL to all encoder blocks yields the best performance.
@@
 # of SBL Normal  Dust  Dark Average
@@
The length of Sequence
We conduct experiments under various $T_{max}$ values. As shown in Table 9, our method shows a significant improvement when extended to video compared to the image-based SCD approach ($T_{max}=1$). However, it is notable that increasing $T_{max}$ does not consistently lead to improved performance; increasing the video length beyond 60 has little to no impact on performance or may even lead to a decline.
@@
The Adaptive Content Threshold
To illustrate the necessity of varying the content threshold according to the sequence length, we conducted experiments across three different sequence lengths. The three sequence lengths were 1, 60, and 30, which are the $T_{max}$ for image-level SCD, our standard $T_{max}$ for video, and the intermediate value, respectively. The fixed threshold values were set to 0.05 and 0.4, approximating the values of ACT when $T_{max}=1$ and $T_{max}=60$, respectively.
As shown in Table 10, when the sequence length is 1, a threshold of 0.05 performs the best, while performance is poor at a threshold of 0.4. Conversely, for sequence lengths of 30 and 60, a threshold of 0.05 results in the lowest performance, while higher thresholds improve performance. Furthermore, the results indicate that the ACT consistently achieves the best performance across all sequence lengths. This shows the validity and effectiveness of ACT and supports the argument that the threshold should be influenced by the sequence length.
@@
$T_{max}$ $\tau$ Normal Dust Dark Average
@@
Table 10: Ablation study on Content Threshold ($\tau$).
Appendix C Qualitative Results
@@
Figure 6: Qualitative Results on ChangeSim Normal subset.
Figure 7: Qualitative Results on ChangeSim Low-illumination and ChangeSim Dusty-air.
Figure 8: Qualitative Results on VL-CMU-CD (Alcantarilla et al. 2018).
Figure 9: Qualitative Results on PCD (Sakurada and Okatani 2015).